OpenStudy (anonymous):

find the system of equations x-y+z=2 x+y+z=4 y+2z=3

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Subtract the second equation from the first.

terenzreignz (terenzreignz):

Or add them :3

OpenStudy (agent0smith):

If you want to do more work, then sure, add them ;)

terenzreignz (terenzreignz):

Crud... really need to pay attention to details >.>

OpenStudy (agent0smith):

hehe :) if you don't like negatives, subtract the first equation from the second.

OpenStudy (anonymous):

okay so umm I see no x y or z's like to help me get started ._.

OpenStudy (campbell_st):

why not a simple substitution the 3rd equation y = 3 - 2z then substitute it into the 1st and 2nd equations and you will have 2 equations in 2 unknowns, x and z. Solve then by elimination

terenzreignz (terenzreignz):

oh.. apologies :) What agent smith meant about subtracting equations is as follows: Say, you have two equations \[\Large \color{blue}{x+y} = \color{red}9\]\[\Large \color{blue}{x-y}=\color{red}{-4}\] Then, say, to subtract the second equation from the first, you just subtract the left-sides, and equate it to the difference of the right-sides... \[\Large (\color{blue}{x+y})-(\color{blue}{x-y})=\color{red}{9}-\color{red}{(-4)}\] \[\Large \color{blue}{2y=\color{red}{13}}\]

OpenStudy (agent0smith):

x-y+z=2 x+y+z=4 y+2z=3 subtract the first from the second to get: x-y+z=2 2y=2 y+2z=3 now it's easy.

terenzreignz (terenzreignz):

Right, I shall steer clear now, lest I confuse the OP :)

OpenStudy (agent0smith):

Because @campbell_st elimination is faster and easier.